# Multimodal Multilingual MOdularized Reciprocal Rank Fusion

### v3

@@ Abstract.
Videos inherently contain multiple modalities, including visual events, text overlays, sounds, and speech, all of which are important for retrieval. However, state-of-the-art multimodal language models like VAST and LanguageBind are built on vision-language models (VLMs), and thus overly prioritize visual signals. Retrieval benchmarks further reinforce this bias by focusing on visual queries and neglecting other modalities. We create a search system MMMORRF that extracts text and features from both visual and audio modalities and integrates them with a novel modality-aware weighted reciprocal rank fusion. MMMORRF is both effective and efficient, demonstrating practicality in searching videos based on users’ information needs instead of visual descriptive queries. We evaluate MMMORRF on MultiVENT 2.0 and TVR, two multimodal benchmarks designed for more targeted information needs, and find that it improves nDCG@20 by 81% over leading multimodal encoders and 37% over single-modality retrieval, demonstrating the value of integrating diverse modalities.
@@
journalyear: 2025conference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–17, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–17, 2025, Padua, Italydoi: XXXisbn: XXXccs: Information systems Video searchccs: Information systems Combination, fusion and federated searchccs: Information systems Multilingual and cross-lingual retrieval
@@ 1. Introduction
Prior research has shown strong results using vision-language models (Chen et al., 2023b; Zhu et al., 2024; Wang et al., 2024) on academic benchmarks like VALOR-32k (Chen et al., 2023a) and MSR-VTT (Xu et al., 2016). However, these datasets are small, English-centric, and rely on descriptive captions that do not reflect real-world search behaviors. Rather than relying on large vision and audio-language models, we introduce a practical and efficient fusion search engine, MMMORRF, that leverages information extracted using mature modality-specific technologies. Figure 1 illustrates our approach, which achieves state-of-the-art results on the MultiVENT 2.0 collection, with more than 80% improvement over the best vision-language system.
Our contributions are three-fold: (1) we present, MMMORRF, a system that combines state-of-the-art components for multilingual, event-centric video retrieval; (2) we introduce a novel modality-aware fusion method for integrating multiple modalities; and (3) we provide a comprehensive comparison with vision-language models, demonstrating the effectiveness of our approach on a more realistic video retrieval task. A video demonstrating MMMORRF is available on YouTube. [^1] https://youtu.be/7jLisTeSjKM
@@ 3. Video Retrieval Pipeline
Visual Content. To capture signals from the visual content, we develop a sub-pipeline that selects, encodes, and searches video frames, illustrated at the top of Figure 1. To balance efficiency and effectiveness, we uniformly sample 16 frames per video. To bridge the modality gap between textual queries and video frames, we use SigLIP (Zhai et al., 2023), a variant of CLIP that replaces the standard contrastive learning softmax normalization with a sigmoid loss function. For efficient retrieval, frame embeddings are indexed using FAISS, a scalable vector search engine (Douze et al., 2024). At search time, queries are encoded using SigLIP’s text encoder and matched against the indexed frame embeddings by computing the maximum query-frame score (MaxFrame) for each video/query pair.
@@
Given the query $q$ and video $d$, the ranks from the two systems are denoted as $r_{text}(q,d)$ and $r_{vision}(q,d)$, respectively, with a smoothing hyperparameter $k$. In our implementation, we set $k=0$, diverging from the typical value of $60$, to emphasize highly ranked results from either modality. This adaptive weighting strategy enables our system to capitalize on the strengths of each modality, enhancing retrieval performance over a diverse video collection.
@@ 4. Experiment Setup
All extracted text from MultiVENT 2.0 is concatenated with their machine translation obtained through Google Translate before PLAID-X indexing, which is a common practice in CLIR (Lawrie et al., 2023a, 2024). For TVR, since the videos are already in English, we do not obtain additional text before indexing.
Following common video retrieval evaluation, we report Recall at 1 and 10 (R@1 and R@10) for both datasets, as well as normalized Discounted Cumulative Gain (nDCG@10) for MultiVENT 2.0 to assess ranked retrieval performance.
@@ 5. Results and Analysis
Pipelines that independently pass OCR and ASR outputs through mature text retrieval systems yield nDCG@10 scores of 0.415 and 0.427 on MultiVENT 2.0, respectively, both outperforming the two vision-language models. In TVR, where the videos are already in English, searching with extracted text still outperforms VAST. Here, searching with PLAID-X does not provide much improvement, indicating that serving these queries can merely rely on lexical matches.

### v4

@@ Abstract.
Videos inherently contain multiple modalities, including visual events, text overlays, sounds, and speech, all of which are important for retrieval. However, state-of-the-art multimodal language models like VAST and LanguageBind are built on vision-language models (VLMs), and thus overly prioritize visual signals. Retrieval benchmarks further reinforce this bias by focusing on visual queries and neglecting other modalities. We create a search system MMMORRF that extracts text and features from both visual and audio modalities and integrates them with a novel modality-aware weighted reciprocal rank fusion. MMMORRF is both effective and efficient, demonstrating practicality in searching videos based on users’ information needs instead of visual descriptive queries. We evaluate MMMORRF on MultiVENT 2.0 and TVR, two multimodal benchmarks designed for more targeted information needs, and find that it improves nDCG@20 by 81% over leading multimodal encoders and 37% over single-modality retrieval.
@@
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730157isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Video searchccs: Information systems Combination, fusion and federated searchccs: Information systems Multilingual and cross-lingual retrieval
@@ 1. Introduction
Prior research has shown strong results using vision-language models (Chen et al., 2023b; Zhu et al., 2024; Wang et al., 2024) on academic benchmarks like VALOR-32k (Chen et al., 2023a) and MSR-VTT (Xu et al., 2016). However, these datasets are small, English-centric, and rely on descriptive captions that do not reflect real-world search behaviors. Rather than relying on large vision and audio-language models, we introduce a practical and efficient fusion search engine, MMMORRF, that leverages information extracted using mature modality-specific technologies. Figure 1 illustrates our approach, which achieves state-of-the-art results on MultiVENT 2.0, with more than 80% improvement over the best vision-language system.
Our contributions are three-fold: (1) we present, MMMORRF, a system that combines state-of-the-art components for multilingual, event-centric video retrieval; (2) we introduce a novel modality-aware fusion method for integrating multiple modalities; and (3) we provide a comprehensive comparison with vision-language models, demonstrating the effectiveness of our approach on a more realistic video retrieval task. A video demonstrating MMMORRF is available on YouTube [^1] https://youtu.be/7jLisTeSjKM, and the implementation is available on GitHub.  [^2] https://github.com/hltcoe/video-retrieval-demo
@@ 3. Video Retrieval Pipeline
Visual Content. To capture signals from the visual content, we develop a sub-pipeline that selects, encodes, and searches video frames, illustrated at the top of Figure 1. To balance efficiency and effectiveness, we uniformly sample 16 frames per video. Through pilot testing on MultiVent 2.0 training set (Kriz et al., 2024), we found that the benefit of sampling more frames saturates dramatically after 16. We also investigated key-frame detection tools, such as PySceneDetect (Castellano, [n.d.]), but found no substantial difference between uniform sampling.
To bridge the modality gap between textual queries and video frames, we use SigLIP (Zhai et al., 2023), a variant of CLIP that replaces the standard contrastive learning softmax normalization with a sigmoid loss function. For efficient retrieval, frame embeddings are indexed using FAISS, a scalable vector search engine (Douze et al., 2024). At search time, queries are encoded using SigLIP’s text encoder and matched against the indexed frame embeddings by computing the maximum query-frame score (MaxFrame) for each video/query pair.
@@
where $\alpha_{d}$ is the coefficient for the combination. Given the query $q$ and video $d$, the ranks from the two systems are denoted as $r_{text}(q,d)$ and $r_{vision}(q,d)$, respectively, with a smoothing hyperparameter $k$. In our implementation, we set $k=0$, diverging from the typical value of $60$, to emphasize highly ranked results from either modality. This adaptive weighting strategy enables our system to capitalize on the strengths of each modality, enhancing retrieval performance over a diverse video collection.
@@ 4. Experiment Setup
All extracted text from MultiVENT 2.0 is concatenated with their machine translation obtained through Google Translate before PLAID-X indexing, which is a common practice in CLIR (Lawrie et al., 2023a, 2024). For TVR, since the videos are already in English, we do not obtain additional text before indexing. Following common video retrieval evaluation, we report Recall at 1 and 10 (R@1 and R@10) for both datasets, as well as normalized Discounted Cumulative Gain (nDCG@10) for MultiVENT 2.0 to assess ranked retrieval performance.
@@
We use eight NVIDIA V100 GPUs for PLAID-X indexing and SigLIP encoding for faster parallel encoding. We ran FAISS without any compression and approximation. Serving queries can be done without a GPU for PLAID-X, but it is necessary for SigLIP. Therefore, at query serving time, we also use a V100 GPU for encoding the queries.
@@ 5. Results and Analysis
Pipelines that independently pass OCR and ASR outputs through mature text retrieval systems yield nDCG@10 scores of 0.415 and 0.427 on MultiVENT 2.0, respectively, both outperforming the two vision-language models. In TVR, where the videos are already in English, searching with extracted text still outperforms VAST. Here, PLAID-X does not provide much improvement, indicating that serving these queries can merely rely on lexical matches.